OpenStudy (anonymous):

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path.The bicyclist heading north is riding 4 km/hour faster than the bicyclist heading south. At 10:30, they are 39 km apart. Find the two bicyclists’ rates. (A)northbound bicyclist = 15 km/h; southbound bicyclist = 11 km/h (B)northbound bicyclist = 17 km/h; southbound bicyclist = 11 km/h (C)northbound bicyclist = 15 km/h; southbound bicyclist = 10 km/h (D)northbound bicyclist = 16 km/h; southbound bicyclist = 12 km/h

OpenStudy (anonymous):

Let x be the speed of the cyclist riding south. Speed of the cyclist reading north = x + 4. 9:00 - 10:30 = 1.5 hours has passed. As Distance = Speed x Time Therefore, Distance travelled by cyclist riding south = (x+4)(1.5) = 1.5x + 6 Distance travelled by cyclist riding north = (x)(1.5) + 1.5x Total distance between them at 10:30 = Distance travelled by cyclist riding south + Distance travelled by cyclist riding north = 1.5x + 6 + 1.5x = 3x + 6 3x + 6 = 39 x = 11 = Southbound Cyclist Northbound Cyclist = 11+4 = 15. Answer: A PS: Always helps to use algebra